OpenStudy (anonymous):

hello... need help please! given: sin A is 4/5 with A in Q2, and cos B = -2/sqrt13 with B in Q3. Find the Following: Cos(A-B) sin2A tan(B/2)

OpenStudy (shiraz14):

cos A = -3/5; sinB = -3/sqrt(13)

OpenStudy (anonymous):

okay, i got that far, but i just cant figure out how to put them in the equations been at this for 2 hours :/

OpenStudy (shiraz14):

You have all the clues from the question and my posts above - just plug them into the equation in my first post in this thread.

OpenStudy (shiraz14):

cos(A-B) = cosAcosB + sinAsinB sin2A = 2sinAcosA tan(B) = [2tan(B/2)]/[1-tan^2(B/2)] (use tanB = sinB/cosB)

OpenStudy (anonymous):

so.... say i did sin2A = 2SinACosA... so it would be 2 * 4/5 *-3/sqrt13 then when you muliply it out, you get -24/5/sqrt13, right?

OpenStudy (anonymous):

when i do cos(A-B), i get 306/325, i think im doing the math wrong, could you show me?

OpenStudy (shiraz14):

In response to your previous question, it should read \[\frac{ -24 }{ 25 }\], not -24/5/sqrt13.

OpenStudy (shiraz14):

In response to your later question, cos(A-B) = (-3/5)(-2/sqrt13) + (4/5){-3/[sqrt(13)]} = 6/(5sqrt13) - 12/(5sqrt13) = -6/(5sqrt13) (Answer)

OpenStudy (anonymous):

could you explain to me how you got -23/25 on the sin2a answer, because i dont how how you got that

OpenStudy (shiraz14):

You used the value of sinB instead of cosA - that was the reason for the miscalculation.

OpenStudy (shiraz14):

cosA = -3/5, not -3/sqrt(13) (which is the value of sinB instead) ...

OpenStudy (anonymous):

so... for 2sinA its 2 * 4/5 * -3/5? i thought you had to use one of each values since it gave you two functions

OpenStudy (shiraz14):

You meant sin(2A). For this, we have: sin(2A) = 2sin(A)cos(A) = 2 * (4/5) * (-3/5) = -24/25 (Answer)

OpenStudy (shiraz14):

Can you try tan(B/2) and show me how you do it? Hint: You'd first need to find tan(B).

OpenStudy (anonymous):

is tan B 3/2?

OpenStudy (shiraz14):

Yes, you're correct. :) Now substitute this value of tanB into the formula in my previous post to find tan(B/2).

OpenStudy (anonymous):

did i enter it in right? i put 2(3/2) over 1 - (3/2)^2

OpenStudy (anonymous):

help? ;(

OpenStudy (anonymous):

can anybody help me? im stuck

OpenStudy (shiraz14):

@GhostShip99 : No, you have to sub. tan(B)=3/2, not tan(B/2) = 3/2 (as you have done). So the expression becomes: 2tan(B/2)/[1-tan^2(B/2)] = 3/2 Then evaluate this expression to find tan(B/2).